# How do Pre-Trained Models Support Software Engineering? An Empirical Study in Hugging Face

### v1

@@ IV-B3 How are cross-activity PTMs applied across multiple SE activities?
To better understand this phenomenon, we selected the PTMs associated with more than one distinct SE activity and represented their pair-wise relationships in Figure 5. The strongest co-occurrence is observed between software implementation and software maintenance, followed by links with software quality assurance and software design. While these numbers highlight absolute overlaps, it is also notable that a substantial proportion of PTMs are reused across activities. For example, 73% of software design PTMs are also applied in software implementation. This suggests that certain PTMs are particularly well-suited for reuse across mid- and late-stage activities in the SDLC, whereas connections to requirements engineering remain relatively sparse.
Figure 5: Co-occurrence graph of cross-activity PTMs.
Finding 2.5: A significant number of PTMs are reused across SE activities, indicating broad applicability. Finding 2.6: Software design PTMs show a particularly strong overlap with software implementation.
IV-B4 Which base models have been most frequently utilized for SE over time?
We analyze the number of SE PTMs developed from specific base models over time using the information provided in the model cards [41]. The “base model” attribute was introduced by HF in November 2023, allowing users to include it in the metadata since then. In this context, users can report whether their PTM is a fine-tuned or adapted version of an original PTM.
Figure 6 presents the top 15 base models with the highest number of PTMs created from them. Established models such as meta-llama/Llama-2-7b-hf[42] have supported 9 PTMs since their introduction, maintaining steady reuse since late 2023. Meanwhile, newer models such as Spestly/Atlas-Flash-7B-Preview[43] have rapidly gained traction, each serving as the base for 36 PTMs within a few months of release. This reflects both the sustained value of proven models and the quick adoption of emerging architectures by the SE community.
Figure 6: Top 15 base models by the number of SE PTMs created.
Finding 2.7: Both established and new base models are widely reused to develop SE PTMs.
IV-B5 What attributes characterize SE PTMs in HF?
@@
SE PTMs can be characterized by a set of core attributes that provide insight into their composition, usage conditions, and performance. To examine how they are documented in practice, we conducted an in-depth analysis of the information provided in the model cards [44] of SE PTMs. Our analysis focused on three main attributes: training dataset, license type, and benchmarks.
Training datasets
@@
High-quality data is essential for effectively training PTMs on domain-specific tasks [45]. To understand how such resources are documented in SE PTMs, we extracted the training dataset information from each model card when available. For the purpose of analysis, we identified the top three training datasets per SE activity based on the number of PTMs that reference each dataset. Notably, if any of these top datasets are reused across multiple SE activities, we explicitly indicate this reuse in Figure 7.
@@
The results reveal clear patterns of dataset reuse across SE tasks. For example, codeparrot/github-code-clean[46] is among the most commonly reused datasets, appearing in 24 PTMs for software implementation and 8 for software quality assurance. This reuse suggests that datasets are not always uniquely tailored to specific SE activities.
@@
In contrast, some datasets show activity-specific use. For instance, jondurbin/airoboros-2.2 [47] appears exclusively in 29 PTMs targeting software implementation, while Tesslate/UIGEN-T1.5-Dataset[48] is used solely in 3 design-related PTMs. Such patterns may reflect more deliberate curation or clearer alignment between dataset content and task requirements.
@@
Figure 7: Top 3 training datasets per SE activity by number of PTMs trained on them.
Finding 2.8: Training datasets are often reused across multiple SE activities. Finding 2.9: No PTMs targeting requirements engineering report associated training datasets.
Licenses
Licensing considerations have become increasingly important in the development and deployment of PTMs, with calls to incorporate license awareness to prevent copyright risks and unintended reuse [49]. In this context, license information provides insights into the legal and practical usage conditions of each PTM. As illustrated in Figure 8, a wide variety of licenses are applied, reflecting different levels of openness, reuse permissions, and restrictions tailored to diverse needs.
Figure 8: License distribution across SE activities.
Permissive licenses such as Apache 2.0 and MIT dominate, particularly in the software implementation phase, where Apache 2.0 licenses cover a striking 857 PTMs and MIT licenses cover 278 PTMs. This dominance aligns with prior works on HF, underscoring the preference for licenses that enable broad use, modification, and redistribution with minimal barriers [50].
More restrictive licenses, such as the Affero General Public License (AGPL-3.0) and the GNU General Public License (GPL-3.0), appear less frequently but are still notable in specific activities (AGPL-3.0 predominantly in software implementation, and GPL-3.0 also appearing in implementation and maintenance phases).
@@
Finding 2.10: Permissive licenses, especially Apache 2.0 and MIT, dominate, facilitating broad adoption and reuse. Finding 2.11: Restrictive licenses are less frequent, but notably present in specific activities.
Benchmarks
PTMs are mastering new benchmarks faster than ever, highlighting the increasing pace of progress in the field [51]. This progress has also reinforced the call for benchmarks that reflect the complexity of real-world development scenarios [52]. Recent surveys of benchmarks for LLMs in SE emphasize the diversity of tasks covered and the need for more standardized evaluation methods [53]. However, there is currently no standardized tag or metadata format in HF that enables consistent reporting of benchmarks across PTMs for straightforward comparison.
Figure 9: Top 20 benchmarks most commonly reported in software implementation PTMs in HF.
After manually reviewing many PTMs during the validation stage (Section III-B5), we decided to automatically search for benchmark tables in model cards. Only PTMs related to software implementation reported benchmarks in tabular form. From these, we identified the top 20 most common benchmarks (Figure 9). The most frequent benchmarks include HumanEval [54] (38 occurrences), Massive Multitask Language Understanding (MMLU) [55] (31), and HellaSwag [56] (31). The lack of a unified standard for reporting benchmarks complicates objective comparisons between PTMs.
Finding 2.12: The absence of standardized benchmark reporting hinders consistent comparison across SE PTMs. Finding 2.13: HumanEval, MMLU, and HellaSwag are the most frequently reported benchmarks among SE PTMs.
V Discussion
A significant gap exists in the documentation of SE PTMs in HF. At the time of the data collection, over 31% of the available PTMs lacked a model card, and only 7.32% included a valid model card that explicitly referenced an SE task. This abscence of documentation reduces transparency, limits informed reuse, and creates obstacles to evaluating the suitability of PTMs for specific SE applications.
@@ V Discussion
Our results align with Hou et al. [16], both identifying a strong focus on code generation tasks, reflecting the broad adoption of generative PTMs in development workflows. As shown in our findings, the number of PTMs associated with software implementation is more than 20 times higher than that of the next most represented SE activity, software maintenance. Our study advances the field by analyzing a substantially larger and more diverse set of over 2,205 PTMs, compared to Hou et al.’s literature-based analysis [16].
@@
Consistent with Yang et al. [17], who focus on code-centric LLMs, we observe this skew toward software implementation. Yet, our work broadens the scope by covering a wider range of SE tasks and phases (identifying 147 distinct SE tasks across five SDLC activities), providing finer granularity and alignment with the full lifecycle.
This imbalance highlights the urgent need for greater research and tooling support for early SDLC activities, such as requirements engineering and software design. The growing interest in PTMs simulating diverse SE roles, such as MetaGPT [57], and the lack of benchmarks for these phases [18], highlight this gap.
Our findings motivate future efforts to extend classification frameworks to additional repositories and develop tools to foster more balanced PTM support across the SDLC, unlocking PTMs’ full potential beyond software implementation.
V-A Implications
The findings of this study provide a practical foundation for both researchers and practitioners seeking to navigate the rapidly growing landscape of SE PTMs. Consider an SE team at a US-based company developing an intelligent code completion tool for Python, integrated into their existing desktop IDE. This team aims to enhance developer productivity through automation while prioritizing environmental sustainability and compliance with open-source licensing policies.
@@ V-A Implications
Our classification supports in identifying and selecting SE PTMs that best align with specific needs, thereby streamlining decision-making processes [58]. Using our catalogue, the team begins by filtering available PTMs for those applicable to software implementation, and they retain only those with explicit support for Python. The next filter selects PTMs that report energy consumption during inference, aligning with their sustainable AI practices. From this subset, they identify PTMs released under permissive licenses, such as MIT. When multiple candidates meet all criteria, the team evaluates them based on quality attributes such as accuracy and benchmark performance. Through this process, the team identifies ashwinR/CodeExplainer[59] as the most suitable, demostrating how our framework can support decision-making in real-world scenarios.
Figure 10 depicts this selection process, emphasizing how our approach guides users to find PTMs best suited to their specific context. Ultimately, this example highlights the framework’s potential to streamline the adoption of SE PTMs, promoting responsible AI development aligned with broader organizational and societal goals.
Figure 10: Selection diagram.
V-B Updatability

### v2

@@ 4.2.2 How are cross-activity PTMs applied across multiple SE activities?
PTMs can be related to more than one SE task, potentially spanning different SE activities (i.e., a PTM can be simultaneously classified into multiple SE activities). To better understand this phenomenon, we selected the PTMs associated with more than one distinct SE activity and represented their pair-wise relationships in Figure 4. The strongest co-occurrence is observed between software implementation and software maintenance, followed by links with software quality assurance and software design.
Although the absolute number of shared PTMs is small relative to the overall corpus of 2,205 SE models, the overlaps are substantial within certain activities. For instance, 22 out of 37 software design PTMs (73%) are also applied in software implementation. This suggests that certain PTMs are particularly well-suited for reuse across mid- and late-stage activities in the SDLC, whereas connections to requirements engineering remain relatively sparse.
Figure 4: Co-occurrence graph of cross-activity PTMs
Finding 2.3: Shared PTMs are few in absolute terms but represent a substantial proportion within certain activities, such as software design.
4.2.3 How are ML tasks related to SE activities?
The relationship between SE activities and ML tasks is illustrated in Figure 5, with the flow indicating the number of PTMs. ML tasks correspond to the current classification used in HF, i.e., a label assigned to each PTM. The figure highlights the prominence of the ML task text generation, which is associated with the highest number of PTMs, reflecting its relevance in supporting code-related tasks. From the SE perspective, this task is most commonly linked to software implementation. However, it also extends across the remaining four SE activities, indicating its versatility.
ML support is uneven across the SDLC. Early phases (requirements engineering and software design) and late phases (software maintenance) demonstrate connections to diverse ML tasks, while middle phases like software design show limited ML task integration, being connected primarily to text generation. In particular, software design is associated with a single ML task (text generation) while the other SE activities are supported by multiple ML tasks. Software quality assurance presents an interesting case with connections to several ML tasks, but notably fewer than software implementation or software maintenance. This highlights a potential growth area for ML applications in testing and validation tasks. Regarding software maintenance, it spans across various ML tasks, including text generation, text classification, and text2text generation. These ML tasks are useful in the context of SE, such as code review tasks.
Figure 5: Association of SE activities with ML tasks
Finding 2.4: The ML task text generation is the most common among SE-related PTMs. Finding 2.5: Unlike other SE activities, software design is related to a single ML task: text generation.
4.3 What attributes characterize SE PTMs in HF? (RQ3)
@@
SE PTMs can be characterized by a set of attributes that provide insights into their properties, performance, and usage conditions. To examine how these attributes are documented in practice, we conducted an in-depth analysis of the information reported in HF model cards huggingfaceModelCards. Our analysis focused on five main attributes: temporal evolution, base model, benchmark scores, training dataset, and license type.
4.3.1 How has the number of SE PTMs evolved over time?
Figure 6 shows a growth in the creation of SE PTMs since 2020, with quarterly data revealing periods of acceleration. Notably, the number of new PTMs developed for software implementation rose from 28 in 2023 Q1 to 142 in 2023 Q2. This upward trend intensified during the last year (from 2024 Q1 to 2024 Q4), peaking at 444 new PTMs in 2024 Q2. In comparison, the current number of PTMs related to software implementation is more than 20 times higher than that of the next most represented SE activity, software maintenance.
In parallel, we observe a steady rise in the number of PTMs associated with requirements engineering, software maintenance, software quality assurance, and software design. Although the absolute number of PTMs for these activities remains much lower than that of software implementation (as shown in Figure 3), the temporal trends indicate a gradual growth across all these SE tasks. This suggests that, while software implementation dominates in volume, other activities are increasingly being addressed by new PTMs.
Figure 6: Number of SE PTMs created quarterly since 2020
Finding 3.1: Since 2020, SE PTMs have grown steadily, with sharp acceleration observed from 2023 onward. Finding 3.2: The number of PTMs addressing SE activities continues to grow, with software implementation maintaining a strong lead.
4.3.2 Which base models have been most frequently utilized for SE over time?
We analyze the number of SE PTMs developed from specific base models over time using the information provided in the model cards huggingfaceModelCardsBaseModel. Figure 7 presents the top 15 base models with the highest number of PTMs created from them.
Figure 7: Top 15 base models by the number of SE PTMs created
Established models such as meta-llama/Llama-2-7b-hf huggingfaceMetallamaLlama27bhfHugging have supported 9 PTMs since their introduction, maintaining steady reuse since late 2023. Meanwhile, newer models such as Spestly/Atlas-Flash-7B-
Preview
huggingfaceSpestlyAtlasFlash7BPreviewHugging have rapidly gained traction, each serving as the base for 36 PTMs within a few months of release. This reflects both the sustained value of proven models and the quick adoption of emerging architectures by the SE community.
Finding 3.3: Both established and new base models are widely reused to develop SE PTMs.
4.3.3 What patterns emerge from the benchmark scores of SE PTMs?
We identified 211 PTMs that included evaluation information (9.57% of the total SE PTMs). Figure 8 illustrates the scores for the ten most frequently reported benchmarks, covering 94.31% of these evaluated PTMs. The distribution reveals that HumanEval, XNLI, and MMLU dominate SE evaluation, accounting for the highest number of reported results, with 1,024, 105, and 103 records, respectively. Notably, despite the large number of PTMs evaluated, most achieve scores below 50% on benchmarks such as BBH, HumanEval, LiveCodeBench, MMLU, MultiPL-HumanEval, RepoQA, and XNLI, indicating that many tasks remain largely unsolved for current PTMs. Among all benchmarks, LiveCodeBench and MultiPL-HumanEval exhibit comparatively lower performances, indicating the difficulty of precise code generation and multi-language evaluation tasks. Conversely, some PTMs reach high scores, up to 94.37% and 90.00% on benchmarks such as XStoryCloze and XCOPA.
Figure 8: Score distribution for the top ten most frequently reported benchmarks in SE PTMs
Finding 3.4: Only 9.57% of SE PTMs report evaluation results, highlighting limited coverage across the ecosystem. Finding 3.5: Most PTMs score below 50% on several common benchmarks, showing many tasks remain largely unsolved. Finding 3.6: XStoryCloze and XCOPA achieve up to 94.37% and 90.00% scores, showing some benchmarks are more tractable for current PTMs.
4.3.4 Which training datasets are used to develop SE PTMs?
@@
For the purpose of analysis, we identified the top three training datasets per SE activity based on the number of PTMs that reference each dataset. Notably, if any of these top datasets are reused across multiple SE activities, we explicitly indicate this reuse in Figure 9.
Figure 9: Top 3 training datasets per SE activity by number of PTMs trained on them
@@
The results reveal clear patterns of dataset reuse across SE tasks. For example, codeparrot/github-code-clean huggingfaceCodeparrotgithubcodecleanDatasets is among the most commonly reused datasets, appearing in 24 PTMs for software implementation and 8 for software quality assurance. This reuse suggests that datasets are not always uniquely tailored to specific SE activities.
In contrast, some datasets show activity-specific use. For instance, jondurbi
n/airoboros-2.2
@@
huggingfaceJondurbinairoboros22Datasets appears exclusively in 29 PTMs targeting software implementation, while Tesslate/UIGEN-T1.5-Dataset huggingfaceTesslateUIGENT15DatasetDatasets is used solely in 3 design-related PTMs. Such patterns may reflect more deliberate curation or clearer alignment between dataset content and task requirements.
@@
Finding 3.7: Training datasets are often reused across multiple SE activities. Finding 3.8: No PTMs targeting requirements engineering report associated training datasets.
4.3.5 What types of licenses are applied to SE PTMs?
As illustrated in Figure 10, a wide variety of licenses are applied, reflecting different levels of reuse permissions and restrictions tailored to diverse needs.
Figure 10: License distribution across SE activities
Permissive licenses such as Apache 2.0 and MIT dominate, particularly in the software implementation phase, where Apache 2.0 licenses cover a striking 857 PTMs and MIT licenses cover 278 PTMs. This dominance aligns with prior works on HF, underscoring the preference for licenses that enable broad use, modification, and redistribution with minimal barriers pepe2024hugging.
More restrictive licenses, such as the Affero General Public License (AGPL-3.0) and the GNU General Public License (GPL-3.0), are less common but notable: AGPL-3.0 predominantly in software implementation, and GPL-3.0 in implementation and maintenance phases.
@@ 4.3.5 What types of licenses are applied to SE PTMs?
Finding 3.9: Permissive licenses, especially Apache 2.0 and MIT, dominate, facilitating broad adoption and reuse. Finding 3.10: Restrictive licenses are less frequent, but notably present in specific activities.
5 Discussion
In this paper, we present a large-scale catalogue of 2,205 SE PTMs available in HF, classified according to a fine-grained list of 147 SE tasks derived from the literature 10.1145/3695988washizaki2024swebok. This classification provides significant insights into SE PTMs, their characteristics, and integration within the SDLC. This section discusses identified documentation and transparency gaps, compares our findings with existing prior studies and large-scale HF datasets, and highlights the practical and research implications of our generated catalogue. Lastly, we detail how our methodology can be updated for future extensions.
5.1 Documentation gaps and transparency in SE PTMs
A significant gap exists in the documentation of SE PTMs in HF. Over 31% of the available PTMs lacked a model card, and only 7.32% included a valid model card that explicitly referenced an SE task. This finding reinforces earlier observations that HF model information is frequently incomplete or inaccurate due to its self-reporting nature of the platform jiang2023empirical. Importantly, this issue is not confined to small or experimental PTMs, but also affects large projects with thousands of downloads stalnaker2025ml.
Prior studies confirm that incomplete documentation is a systemic problem across HF. horwitz2025charting, in their analysis of 63,000 models from HF, found that only 8% of them outlined license information, 13.4% reported accuracy details, and 30.2% specified their base model. Similarly, suryani2025model, in a broader-scale study covering the entire HF, reported that nearly one-third of model cards include licensing information. In our SE-focused dataset, only 9.57% of PTMs report evaluation results, often using different metrics and benchmarks, further complicating comparative assessment. Licensing patterns among SE PTMs also reflect broader trends across the HF platform. As reported by stalnaker2025ml, 62.5% of all HF models are released under open-source licenses, with apache-2.0 and mit emerging as the most common choices, consistent with our findings for SE PTMs. Notably, in our data, the “other” license category ranks as the third most frequent among SE PTMs, while in the overall HF corpus it ranks fifth stalnaker2025ml. As discussed by stalnaker2025ml, this category often appears when PTMs were uploaded before the correct license was included in HF’s official list or when model owners are unfamiliar with available licensing options.
These findings align with ecosystem-level trends identified by laufer2025anatomy, who found that model cards within families tend to shorten and become increasingly standardized through templates or automatically generated text laufer2025anatomy, mirroring our observation of model card missing information and similar documentation across PTMs. This absence of documentation reduces transparency, limits informed reuse, and creates obstacles to evaluating the suitability of PTMs for specific SE applications.
5.2 Comparison with prior studies
@@
Our results align with 10.1145/3695988, both identifying a strong focus on code generation tasks, reflecting the broad adoption of generative PTMs in development workflows. As shown in our findings, the number of PTMs associated with software implementation is more than 20 times higher than that of the next most represented SE activity, software maintenance. Our study advances the field by analyzing a substantially larger and more diverse set of over 2,205 PTMs, compared to the literature-based analysis by 10.1145/3695988.
@@
Consistent with yang2024ecosystemlargelanguagemodels, who focus on code-centric LLMs, we observe this skew toward software implementation. Yet, our work broadens the scope by covering a wider range of SE tasks and phases (identifying 147 distinct SE tasks across five SDLC activities), providing finer granularity and alignment with the full lifecycle. This imbalance highlights the urgent need for greater research and tooling support for early SDLC activities, such as requirements engineering and software design. The growing interest in PTMs simulating diverse SE roles, such as MetaGPT hong2023metagpt, and the lack of benchmarks for these phases wang2025software, highlight this gap.
ferino2025novice focused on LLM adoption among novice developers, highlighting that the most addressed tasks belonged to the software development activity and software maintenance. At the same time, they also identified challenges (i.e., losing learning opportunities) and suggested best practices (e.g., prompt engineering). Nevertheless, their study concentrates only on LLMs, with a focus on novice developers’ practices.
In terms of SDLC representation, Figure 11 illustrates the number of models identified by each of the three aforementioned studies (10.1145/3695988; yang2024ecosystemlargelanguagemodels; ferino2025novice) and ours. All four studies show that software implementation is the most dominant SE activity in terms of model coverage, a trend that is most pronounced in our study, with 2,004 PTMs.
Figure 11: Distribution of models per SE activity among related works that are aligned to the SDLC
Regarding requirements engineering, we have catalogued a similar amount of resources as 10.1145/3695988, but significantly less than ferino2025novice. Notably, ferino2025novice grouped multiple SE tasks into broader SE activities, which, for instance, included 20 studies employing LLMs for benchmarking as part of the combined activity “Requirement Design & Software Design”.
For software design, all previous studies analyzed a similar number of models (4, 5, and 7), while we found up to four times more PTMs in HF for such activity (30 PTMs), which may be explained by our broader scope and the inclusion of more recent PTMs.
For the case of software quality assurance, our study identified a considerable jump, cataloging 129 models, significantly surpassing the next highest count of 99 by 10.1145/3695988, and exceeding the counts of 25 and 24 from wang2025software and ferino2025novice, respectively. This suggests greater activity for software quality assurance that previous studies may have missed.
Regarding software maintenance, our study again registered the highest count, with 148 models, compared to 66 for 10.1145/3695988, 24 for wang2025software, and 73 for ferino2025novice This indicates that, while still far behind implementation, maintenance is a growing area for SE PTMs.
The overwhelming prevalence of software implementation PTMs (2,704 models) compared to early-stage activities like requirements engineering (27 models) warrants specific discussion. This disparity is likely driven by the abundance of structured code datasets and mature execution-based benchmarks which facilitate development.
5.3 Overlap with large HF datasets
We have also compared how our catalogue overlaps with two large, generic HF-based datasets to contextualize its coverage. HFCommunity 10123660 aggregates data from the HF Hub, while PeaTMOSS jiang2024peatmoss comprises metadata for 281,638 PTMs and detailed snapshots for those with over 50 monthly downloads (14,296 PTMs), and links PTMs with 28,575 open-source software repositories from GitHub that use them.
We found that $1,338$ ($43.48\%$) and 382382 ($12.38\%$) of our SE PTMs appear in HFCommunity and PeaTMOSS, respectively. This overlap ($0.2\%$ of those collections), arises from the recency of our catalogue, as HFCommunity has been static since October 2024 and PeaTMOSS since August 2023, and from the inclusivity of our catalogue, which retains specialized and emerging PTMs regardless of their popularity.
Table 7 quantifies this temporal gap by illustrating the number of PTMs in our study released after the cutoff dates of the comparison datasets. As both have been static for a long period, they miss a significant portion of the recent surge in SE models. For instance, we identified 2,332 software implementation models and 142 quality assurance models that were published after the PeaTMOSS snapshot. Even compared to the more recent HFCommunity, our catalogue captures 921 additional implementation models released after October 2024. This demonstrates the rapid evolution of the ecosystem. However, these partial intersections create opportunities for richer analysis, since combining our catalogue with broader datasets can reveal ecosystem-level insights that remain hidden when these resources are considered in isolation.
Table 7: Number of SE PTMs in our catalogue published after the data collection cutoff dates of PeaTMOSS and HFCommunity
SE Activity
After PeaTMOSS
(Aug 2023)
After HFCommunity
(Oct 2024)
Requirements engineering 18 8
Software design 36 24
Software implementation 2,332 921
Software maintenance 130 59
Software quality assurance 142 63
5.4 Implications
The findings of this study provide a practical foundation for both researchers and practitioners seeking to navigate the rapidly growing landscape of SE PTMs. By aligning PTMs with SDLC, our catalogue supports informed selection based on project needs, such as license type or reported evaluation.
@@ 5.4 Implications
From a practitioner’s perspective, our framework facilitates the responsible selection and reuse of PTMs in software projects. Consider an SE team at a US-based company developing a code completion tool for Python, integrated into their existing desktop IDE. This team aims to enhance developer productivity through automation while prioritizing environmental sustainability and compliance with open-source licensing policies. Our catalogue supports selecting SE PTMs that best align with specific needs, thereby streamlining decision-making processes ben2025should. As characterized by zhouunifying, a combination of semantic and metric-based selection criteria can be set. In this case, the team queries the database for PTMs catalogued under software implementation and refines the results to those explicitly supporting Python. They then restrict the selection to PTMs that report energy consumption, aligning with their sustainable Artificial Intelligence (AI) practices. From this subset, they identify PTMs released under permissive licenses, such as MIT. When multiple candidates meet all criteria, the team evaluates them based on quality attributes such as accuracy and benchmark performance. Through this process, the team identifies ashwinR/CodeExplainer huggingfaceAshwinRCodeExplainerHugging as the most suitable, demonstrating how our framework can support decision-making in real-world scenarios. Figure 12 depicts this selection process, emphasizing how our approach guides users to find PTMs best suited to their specific context. Ultimately, this example highlights the framework’s potential to streamline the adoption of SE PTMs, promoting responsible AI development aligned with broader organizational and societal goals.
Figure 12: Selection diagram
The findings of this study have also motivated the ongoing development of MLAssetSelection, a web-application that operationalizes our cataloguing pipeline to assist in the sampling and selection of SE PTMs mlassetselection. As illustrated in Figure 13, the tool provides an interactive interface for exploring the catalogue. Users can filter PTMs by identifier, scope (e.g., SE activities and tasks, ML tasks), metadata (e.g., licenses, libraries, natural languages, creation date, size categories: bytes for models and rows for datasets), popularity (e.g., downloads, likes), activity (e.g., commits, contributors), and model-specific attributes (e.g., training dataset, region, inference providers, evaluation).
Figure 13: MLAssetSelection overview
From a research perspective, our findings reveal important gaps and opportunities in the current SE PTM ecosystem. Certain phases of the SDLC remain underrepresented, particularly requirements engineering and software design, where few PTMs explicitly target these activities. Beyond identifying such gaps, our catalogue provides a foundation for researchers to delve deeper into the challenges uncovered in this study and to conduct empirical studies using the collected SE PTMs.
5.5 Updatability